OpenStudy (anonymous):

Helpful steps please, A vibrating wire has a vertical displacement given by y=8(cos (2 pi x) cos (bt) + sin (2 pi x) sin (bt) ) assume b and t are constant. write y as a cosine function of x

OpenStudy (perl):

do you want to remove the sine from the function

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

Recall the identity, \[\cos(x-y)=\cos x\cos y+\sin x\sin y\]

OpenStudy (anonymous):

would it then be y=8cos(2 pi x-bt)?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thanks